# DemoRank: Selecting Effective Demonstrations for Large Language Models in Ranking Task

### v1

@@
Abstract
Recently, there has been increasing interest in applying large language models (LLMs) as zero-shot passage rankers. However, few studies have explored how to select appropriate in-context demonstrations for the passage ranking task, which is the focus of this paper. Previous studies mainly apply a demonstration retriever to retrieve demonstrations and use top-$k$ demonstrations for in-context learning (ICL). Although effective, this approach overlooks the dependencies between demonstrations, leading to inferior performance of few-shot ICL in the passage ranking task. In this paper, we formulate the demonstration selection as a retrieve-then-rerank process and introduce the DemoRank framework. In this framework, we first use LLM feedback to train a demonstration retriever and construct a novel dependency-aware training samples to train a demonstration reranker to improve few-shot ICL. The construction of such training samples not only considers demonstration dependencies but also performs in an efficient way. Extensive experiments demonstrate DemoRank’s effectiveness in in-domain scenarios and strong generalization to out-of-domain scenarios. Our codes are available at https://github.com/8421BCD/DemoRank.
DemoRank: Selecting Effective Demonstrations for Large Language Models in Ranking Task
Wenhan Liu, Yutao Zhu,  and Zhicheng Dou Gaoling School of Artificial Intelligence, Renmin University of China lwh@ruc.edu.cn, yutaozhu94@gmail.com, dou@ruc.edu.cn
1 Introduction
@@
Large language models (LLM) have demonstrated remarkable performance across a spectrum of natural language processing (NLP) tasks. Recently, there has been significant interest in using LLMs for passage ranking tasks Zhuang et al. (2023a); Sun et al. (2023); Qin et al. (2023). A typical approach is relevance generation, which judges the relevance of a query-passage pair in a pointwise manner. This method prompts LLMs to assess the relevance of a passage to a query by generating responses such as “Yes” or “No”. The relevance score is then computed based on the log-likelihood of these responses. This approach has been demonstrated to be effective in previous studies Zhuang et al. (2023a); Liang et al. (2022).
@@
Figure 1: Compared with choosing top-2 demonstrations ($z_{1}$ and $z_{2}$), the combination of $z_{1}$ and $z_{5}$ provides richer and more diverse query-passage relationships, thus yielding better relevance assessment.
@@
In-context learning (ICL) has been proved as an emergent ability of LLMs Wei et al. (2022), enabling them to adapt to specific tasks through several task demonstrations (i.e., input-output examples). Many studies have investigated the optimal selection of demonstrations for NLP tasks Lu et al. (2022); Zhang et al. (2022); Li et al. (2023); Wang et al. (2023); Xu et al. (2024), highlighting the importance of tailored demonstrations in achieving high performance. However, the application of ICL to passage ranking tasks has not been extensively studied. Given the complex nature of passage ranking, ICL presents a challenging yet promising opportunity to enhance LLMs’ performance. Consequently, this study aims to develop effective demonstration selection strategies to optimize the application of ICL in passage ranking.
@@
A widely-used and effective approach for demonstration selection is training a demonstration retriever using LLM’ feedback Wang et al. (2023); Rubin et al. (2022); Li et al. (2023); Cheng et al. (2023); Scarlatos and Lan (2023); Luo et al. (2023). This approach first utilizes an LLM to score some demonstration candidates based on LLM’s likelihood of producing the correct output given each candidate and the input, and choose positive and negative candidates based on scores for retriever training. Following this technique line, we propose to train a demonstration retriever based on LLM’s feedback tailored for passage ranking task.
@@
In the inference stage, a common practice Wang et al. (2023) is to use the trained retriever to obtain a list of demonstrations and concatenate the top-retrieved ones together in the prompt for ICL. Despite its effectiveness in NLP tasks, directly extending it into the passage ranking task may result in sub-optimal performance. The main challenge lies in the complex nature of the query-passage relationship in passage ranking, which may require a combination of multiple demonstrations to provide effective information for understanding such a relationship. Figure 1 shows an example of such a problem. When selecting 2-shot demonstrations for the current input (a relevant query-passage pair), existing methods Wang et al. (2023); Rubin et al. (2022) will choose the top-2 demonstrations ($z_{1}$ and $z_{2}$) returned by the retriever. However, we deem that combining $z_{1}$ and $z_{5}$ is more suitable for this case. This is because $z_{1}$ and $z_{5}$ have more distinct queries and opposite outputs (relevance label), which provide LLM with richer and more diverse query-passage relationship signals, thus contributing more to the relevance assessment. This example shows the insufficiency of pure relevance-based demonstration selection in the few-shot LLM-based passage ranking task. In this paper, we transform the problem of selecting the optimal $k$-shot demonstrations from initially retrieved $n$ demonstrations into a demonstration ranking problem and propose to use LLM’s feedback to train a novel dependency-aware demonstration reranker, making the top-ranked ones more suitable in the few-shot ICL for passage ranking.
@@
Nevertheless, training such a reranker is a very challenging task. As previously mentioned, it is unreasonable to use LLM’s feedback on each individual demonstration for training a reranker designed for $k$-shot selection, because demonstrations can influence each other. Additionally, constructing the ground truth ranking of a reranker tailored for $k$-shot selection requires finding the optimal $k$-shot permutation from the retrieved $n$ demonstrations. Theoretically, this requires using LLM to score total $\frac{n!}{(n-k)!}$ demonstration permutations, which is highly time-consuming and impractical. To overcome these challenges, we propose to construct a kind of dependency-aware training samples (a list of demonstrations with ranking labels) for reranker training. Specifically, given a retrieved demonstration set, we greedily select demonstrations from the set and annotate them with different ranking labels (from highest to lowest). Each time, the demonstration that maximizes the LLM’s feedback when concatenated with the already selected ones is chosen. This process not only considers the dependencies between current demonstration and previously selected ones, but also greatly reduces the number of LLM inferences.
@@
To this end, we propose DemoRank, a Demonstration selection framework for passage Ranking, using a two-stage “retrieve-then-rerank” strategy. In this framework, we first train a demonstration retriever DRetriever based on LLM’s feedback for the ranking task. Then, we introduce a dependency-aware demonstration reranker DReranker to rerank the retrieved demonstrations. To address the challenges of its training, we propose a method to construct dependency-aware training samples that not only incorporates demonstration dependency but is also time-efficient.
@@
Experiments on a series of ranking datasets prove the effectiveness of DemoRank, especially in few-shot ICL. Further analysis also demonstrates the contribution of each proposed component and DemoRank’s strong ability under different scenarios, including limited training data, different demonstration numbers, unseen datasets, etc.
The main contributions of our paper are summarized as follows: (1) To the best of our knowledge, we are the first to comprehensively discuss effective demonstration selection in passage ranking and propose DemoRank framework. (2) We propose a novel dependency-aware demonstration reranker and design a rational and efficient method for constructing its training data. (3) Besides in-domain performance, further experiments also demonstrate DemoRank’s generalization on unseen datasets.
2 Related Work
2.1 LLM for Passage Ranking
@@
With the development of large language models (LLMs) in information retrieval Zhu et al. (2023), there have been many studies exploring how to utilize LLMs for the passage ranking task. In general, these studies can be divided into three categories: pointwise Liang et al. (2022); Sachan et al. (2022), pairwise Qin et al. (2023), and listwise methods Sun et al. (2023); Ma et al. (2023). Pointwise methods assess the relevance between a query and a single passage. A typical approach is relevance generation Liang et al. (2022); Zhuang et al. (2023a), which provides LLM with a query-passage pair and instructs it to output “Yes” if the passage is relevant to the query or “No” if not. The relevance score can be calculated based on the generation probability of the token “Yes”. Another approach of pointwise methods is query generation Sachan et al. (2022); Zhuang et al. (2023b), which calculates relevance score based on the log-likelihood of generating the query based on the passage. Pairwise methods compare two passages at a time and determine their relative relevance to a query, and listwise methods directly rank a passage list.
@@
Despite promising results, these studies only focus on the zero-shot scenarios, with less emphasis on how to select effective demonstrations in few-shot scenarios. Manually written or rule-based selection Drozdov et al. (2023) is inflexible for ranking tasks. In this paper, we explore more effective demonstration selection approaches for ranking tasks. Previous studies Zhu et al. (2024) have revealed that relevance generation of the pointwise method is the most suitable method for passage ranking on open-source LLMs compared with other methods. Thus, we intend to use the relevance generation approach for passage ranking in this paper.
2.2 Demonstration Retrieval
@@
A widely used demonstration selection approach is demonstration retrieval. Prior studies have explored using different retrievers for demonstration retrieval, which can be divided into two categories. One is utilizing off-the-shelf retrievers such as BM25 Agrawal et al. (2023) or dense retriever Liu et al. (2022). The other is to train a demonstration retriever using task-specific signals. For example, Rubin et al. (2022) propose to distill the LLM’s feedback to a dense retriever EPR for the semantic parsing task. Li et al. (2023) and Wang et al. (2023) propose to train the retriever iteratively on various NLP tasks. However, a common issue with these methods is that they directly choose the top-retrieved demonstrations, which may include redundant information and contribute little to the LLM’s understanding of relevance. In this paper, we take the demonstration dependencies into account and introduce a framework that first retrieves a list of demonstrations and then reranks in a dependency-aware manner, better aligning with the few-shot ICL in the ranking task.
3 Preliminaries
3.1 Relevance Generation for Ranking Task
@@
Passage ranking aims to rank a list of retrieved passages based on their relevance to a query. Formally, given a query $q$ and a passage list $\left[p_{1},\dots,p_{n}\right]$, our task is to compute a relevance score $S(q,p_{i})$ for each passage. In the LLM-based relevance generation methods Liang et al. (2022); Zhuang et al. (2023a), an LLM is provided with a prompt consisting of a query and a passage, and instructed to output a binary label “Yes” or “No” to indicate whether the passage is relevant to the query or not. Then a softmax function is applied to the logits of tokens “Yes” and “No”, and the probability of the token “Yes” is used as the relevance score:
$Rs(q,p_{i})=\text{Pr}(\text{``Yes''}|T,q,p_{i}),$ (1)
@@
3.2 In-context Learning in Ranking Task
@@
In-context learning is a technique that inserts a few demonstrations into the prompt to help LLMs perform a task without updating parameters. In relevance generation task, given $k$ in-context demonstrations $\{z_{i}\}_{i=1}^{k}$, where $z_{i}=(\hat{q},\hat{p},\hat{y})$ is a triple consisting of a query, a passage and a binary output (“Yes” or “No”) indicating the relevance label, the relevance score $Rs(q,p_{i})$ could be calculated by:
$Rs(q,p_{i})=\text{Pr}(\text{``Yes''}|T,\{z_{i}\}_{i=1}^{k},q,p_{i}),$ (2)
@@
where $T$ is the task description, which is used in ICL to help LLMs understand the task Zhu et al. (2024); Li et al. (2023).
@@
Figure 2: An overview of our proposed framework DemoRank. DemoRank comprises two main components: DRetriever and DReranker. We train the DRetriever using demonstration candidates scored by LLM and construct a kind of dependency-aware training samples to train the DReranker. During inference, a retrieve-then-rerank pipeline is performed and the top-$k$ reranked ones are used for ICL.
4 The DemoRank Framework
@@
As shown in Figure 2, our DemoRank framework follows a process of demonstration retrieval followed by dependency-aware reranking. The demonstration retriever DRetriever is trained using the demonstration candidates scored by LLM and the demonstration reranker DReranker is trained based on our constructed dependency-aware training samples. In this section, we elaborate on our demonstration pool construction, the pipeline of training, and inference.
4.1 Demonstration Pool Construction.
@@
Given a passage ranking dataset (e.g., MS MARCO Nguyen et al. (2016)), we use its training set to construct our demonstration pool $\mathcal{P}$. For each query in the training set, we construct positive and negative demonstrations by pairing the query with its relevant and irrelevant passages respectively. To maintain the output label balance in the demonstration pool $\mathcal{P}$, the number of negative demonstrations of each query is set equal to its positive demonstrations.
4.2 Demonstration Retriever DRetriever

### v2

@@
Abstract.
Recently, there has been increasing interest in applying large language models (LLMs) as zero-shot passage rankers. However, few studies have explored how to select appropriate in-context demonstrations for the passage ranking task, which is the focus of this paper. Previous studies mainly use LLM’s feedback to train a retriever for demonstration selection. These studies apply the LLM to score each demonstration independently, which ignores the dependencies between demonstrations (especially important in ranking task), leading to inferior performance of top-$k$ retrieved demonstrations. To mitigate this issue, we introduce a demonstration reranker to rerank the retrieved demonstrations so that top-$k$ ranked ones are more suitable for ICL. However, generating training data for such reranker is quite challenging. On the one hand, different from demonstration retriever, the training samples of reranker need to incorporate demonstration dependencies. On the other hand, obtaining the gold ranking from the retrieved demonstrations is an NP-hard problem, which is hard to implement. To overcome these challenges, we propose a method to approximate the optimal demonstration list iteratively and utilize LLM to score demonstration lists of varying lengths. By doing so, the search space is greatly reduced and demonstration dependencies are considered. Based on these scored demonstration lists, we further design a list-pairwise training approach which compares a pair of lists that only differ in the last demonstration, to teach the reranker how to select the next demonstration given a previous sequence. In this paper, we propose a demonstration selection framework DemoRank for ranking task and conduct extensive experiments to prove its strong ability.
Passage Ranking, Large Language Model, In-context Learning
1. Introduction
@@
Large language models (LLM) have demonstrated remarkable performance across a spectrum of natural language processing (NLP) tasks. Recently, there has been significant interest in using LLMs for passage ranking tasks (Zhuang et al., 2023b; Sun et al., 2023; Qin et al., 2023). A typical approach is relevance generation, which judges the relevance of a query-passage pair in a pointwise manner. This method prompts LLMs to assess the relevance of a passage to a query by generating responses such as “Yes” or “No”. The relevance score is then computed based on the log-likelihood of these responses. This approach has been demonstrated to be effective in previous studies (Zhuang et al., 2023b; Liang et al., 2022).
@@
Figure 1. Compared with choosing top-2 demonstrations ($z_{1}$ and $z_{2}$), the combination of $z_{1}$ and $z_{5}$ provides richer and more diverse query-passage relationships, thus yielding better relevance assessment.
@@
In-context learning (ICL) has been proved as an emergent ability of LLMs (Wei et al., 2022), enabling them to adapt to specific tasks through several task demonstrations (i.e., input-output examples). This mechanism facilitates the LLMs’ comprehension of both the objective and the expected output format of the task, thereby enhancing performance. Many studies have investigated the optimal selection of demonstrations for NLP tasks (Lu et al., 2022; Zhang et al., 2022; Li et al., 2023; Wang et al., 2023; Xu et al., 2024), highlighting the importance of tailored demonstrations in achieving high performance. However, the application of ICL to passage ranking tasks has not been extensively studied. Given the complex nature of passage ranking, ICL presents a challenging yet promising opportunity to enhance LLMs’ performance. Consequently, this study aims to develop effective demonstration selection strategies to optimize the application of ICL in passage ranking.
@@
A widely-used and effective approach for demonstration selection is training a demonstration retriever using LLM’s feedback (Wang et al., 2023; Rubin et al., 2022; Li et al., 2023; Cheng et al., 2023; Scarlatos and Lan, 2023; Luo et al., 2023). This approach first utilizes an LLM to score some demonstration candidates based on LLM’s likelihood of producing the correct output given each candidate and the input, and choose positive and negative candidates based on scores for retriever training. Following this technique line, we propose to train a demonstration retriever tailored for passage ranking task based on LLM’s feedback.
@@
In the inference stage, a common practice (Wang et al., 2023) is to use the trained retriever to retrieve a list of demonstrations and concatenate the top-retrieved ones together in the prompt for ICL. Despite its effectiveness in NLP tasks, directly extending it into the passage ranking task may result in sub-optimal performance. The main challenge lies in the complex nature of the query-passage relationship in passage ranking, which may require a combination of multiple demonstrations to provide effective information for understanding such a relationship. Figure 1 shows an example of such problem. When selecting a 2-shot demonstration for the current input (a relevant query-passage pair), existing methods (Wang et al., 2023; Rubin et al., 2022) will choose the top-2 demonstrations ($z_{1}$ and $z_{2}$) returned by the retriever. However, we deem that combining $z_{1}$ and $z_{5}$ is more suitable for this case. This is because $z_{1}$ and $z_{5}$ have more distinct queries and opposite outputs (relevance label), which provide LLM with richer and more diverse query-passage relationship signals, thus contributing more to the relevance assessment. This example shows the insufficiency of pure relevance-based demonstration selection in the few-shot LLM-based passage ranking task. In this paper, we transform the problem of selecting the optimal $k$-shot demonstration from retrieved $n$ demonstrations into a demonstration reranking problem and propose to train a dependency-aware demonstration reranker that iteratively selects the next demonstration given the previous demonstration sequence.
@@
Nevertheless, training such a reranker is a very challenging task. As previously mentioned, it is unreasonable to use LLM’s feedback on each individual demonstration for training a reranker designed for $k$-shot selection, because demonstrations can influence each other. Additionally, constructing the ground truth ranking of a reranker tailored for $k$-shot selection requires finding the optimal $k$-shot permutation from the retrieved $n$ demonstrations, which is an NP-hard problem. Theoretically, this requires using LLM to score total $\frac{n!}{(n-k)!}$ demonstration permutations, which is highly time-consuming and impractical. To overcome these challenges, we propose an efficient approach to construct a kind of dependency-aware training samples. Specifically, our approach approximates the optimal demonstration list by iteratively selecting demonstrations from retrieved ones, which significantly reduces the search space. In each iteration, we use an LLM to score a set of demonstration lists that differ only in the last demonstration (e.g., ${[z_{3},z_{1},z_{2}],[z_{3},z_{1},z_{4}],[z_{3},z_{1},z_{5}],\dots}$). Taking the whole list as input, the LLM score considers the dependencies among demonstrations. After constructing the training samples, we design a novel list-pairwise training method which compares a pair of demonstration lists (e.g., $([z_{3},z_{1},z_{2}],[z_{3},z_{1},z_{4}])$), so that the reranker learns to select the next demonstration given a previous sequence.
@@
To this end, we propose DemoRank, a Demonstration selection framework for passage Ranking, using a two-stage “retrieve-then-rerank” strategy. In this framework, we first introduce a demonstration retriever DRetriever trained using demonstrations individually scored by LLM. Then, we introduce a dependency-aware demonstration reranker DReranker to iteratively select demonstrations, so as to obtain a reranked demonstration list. To address the challenges of its training, we propose a method to construct dependency-aware training samples that not only incorporates demonstration dependency but is also time-efficient. We also propose a list-pairwise approach for DReranker training, which focuses on comparing a pair of demonstration lists.
@@
Experiments on a series of ranking datasets prove the effectiveness of DemoRank, especially in few-shot ICL. Further analysis also demonstrates the contribution of each proposed component and DemoRank’s strong ability under different scenarios, including limited training data, different demonstration numbers, unseen datasets, different LLM rankers, etc.
The main contributions of our paper are summarized as follows:
(1) To the best of our knowledge, we are the first to comprehensively discuss effective demonstration selection in passage ranking and propose DemoRank which consists of a demonstration retriever and a dependency-aware demonstration reranker.
(2) We propose an efficient method for constructing dependency-aware training samples for demonstration reranker.
(3) Based on these training samples, we design a list-pairwise training approach to teach the reranker to select the next demonstration based on some contexts.
2. Related Work
2.1. LLM for Passage Ranking
@@
With the development of large language models (LLMs) in information retrieval (Zhu et al., 2023), there have been many studies exploring how to utilize LLMs for the passage ranking task. In general, these studies can be divided into three categories: pointwise (Liang et al., 2022; Sachan et al., 2022), pairwise (Qin et al., 2023), and listwise methods (Sun et al., 2023). Pointwise methods assess the relevance between a query and a single passage. A typical approach is relevance generation (Liang et al., 2022; Zhuang et al., 2023b), which provides LLM with a query-passage pair and instructs it to output “Yes” if the passage is relevant to the query or “No” if not. The relevance score can be calculated based on the generation probability of the token “Yes”. Another approach of pointwise methods is query generation (Sachan et al., 2022; Zhuang et al., 2023a), which calculates relevance score based on the log-likelihood of generating the query based on the passage. Pairwise methods compare two passages at a time and determine their relative relevance to a query, and listwise methods directly rank a passage list.
@@
Despite promising results, these studies only focus on the zero-shot scenarios, with less emphasis on how to select effective few-shot demonstrations. Manually written or rule-based selection (Drozdov et al., 2023) is inflexible for ranking tasks. In this paper, we explore more effective few-shot demonstrations selection approaches for ranking tasks. Previous studies (Zhu et al., 2024) have revealed that relevance generation of the pointwise method is the most suitable method for passage ranking on open-source LLMs compared with other methods. Thus, we intend to use the relevance generation approach for passage ranking in this paper.
2.2. Demonstration Retrieval
@@
A widely used demonstration selection approach is demonstration retrieval. Prior studies have explored using different retrievers for demonstration retrieval, which can be divided into two categories. One is utilizing off-the-shelf retrievers. For example, Agrawal et al. (2023) propose to use BM25 for demonstration retrieval in machine translation. Liu et al. (2022) propose to use a kNN-based retriever to retrieve demonstrations semantically similar to the test input. The other is to train a demonstration retriever using task-specific signals. For example, Rubin et al. (2022) propose to distill the LLM’s feedback to a dense retriever EPR for the semantic parsing task.  Cheng et al. (2023) focuses on training a demonstration retriever that improves LLMs’ performance in the cross-task and cross-model scenarios. Li et al. (2023) and Wang et al. (2023) propose to train the retriever iteratively on various NLP tasks. However, a common issue with these methods is that they directly choose the top-retrieved demonstrations, which may include redundant information and contribute little to the LLM’s understanding of relevance. In this paper, we take the demonstration dependencies into account and introduce a framework that first retrieves a list of demonstrations and then reranks in a dependency-aware manner, better aligning with the few-shot ICL in the ranking task.
3. Preliminaries
3.1. Relevance Generation for Ranking Task
@@
Passage ranking aims to rank a list of retrieved passages based on their relevance to a query. Formally, given a query $q$ and a passage list $\left[p_{1},\dots,p_{n}\right]$, our task is to compute a relevance score $S(q,p_{i})$ for each passage. In the LLM-based relevance generation methods (Liang et al., 2022; Zhuang et al., 2023b), an LLM is provided with a prompt consisting of a query and a passage, and instructed to output a binary label “Yes” or “No” to indicate whether the passage is relevant to the query or not. Then a softmax function is applied to the logits of tokens “Yes” and “No”, and the probability of the token “Yes” is used as the relevance score:
(1) $Rs(q,p_{i})=\text{Pr}(\text{``Yes''}|T,q,p_{i}),$
@@
3.2. In-context Learning in Ranking Task
@@
In-context learning is a technique that inserts a few demonstrations into the prompt to help LLMs perform a task without updating parameters. In relevance generation task, given $k$ in-context demonstrations $\{z_{i}\}_{i=1}^{k}$, where $z_{i}=(\hat{q},\hat{p},\hat{y})$ is a triple consisting of a query, a passage and a binary output (“Yes” or “No”) indicating the relevance label, the relevance score $Rs(q,p_{i})$ is calculated by:
(2) $Rs(q,p_{i})=\text{Pr}(\text{``Yes''}|T,\{z_{i}\}_{i=1}^{k},q,p_{i}),$
@@
where $T$ is the task description, which is used in ICL to help LLMs understand the task (Zhu et al., 2024; Li et al., 2023).
@@
Figure 2. An overview of our proposed framework DemoRank. It comprises two main components: DRetriever and DReranker. We first train the DRetriever using demonstration candidates individually scored by LLM. Then, We iteratively construct a set of dependency-aware training samples (each sample is a ranking of a $k$-shot demonstration list) and use a list-pairwise training approach to optimize the DReranker. During inference, we first use DRetriever to obtain a list of demonstrations and then apply DReranker to iteratively select demonstrations to obtain the top-$k$ reranked demonstrations for ICL.
4. The DemoRank Framework
@@
As shown in Figure 2, our DemoRank framework follows a process of demonstration retrieval and dependency-aware reranking. We construct the training data for demonstration retriever DRetriever by individually scoring each demonstration candidate. As for the demonstration reranker DReranker, as we mentioned in Section 1, it is unreasonable to individually score each demonstration and impractical to obtain the optimal demonstration ranking. To overcome these challenges, we construct a kind of dependency-aware training samples in an efficient way for DReranker training. Based on these training samples, we further propose a list-pairwise training approach for model optimization. In this section, we will elaborate on our demonstration pool construction, the pipeline of model training, and inference.
4.1. Demonstration Pool Construction.
@@
Given a passage ranking dataset (e.g., MS MARCO (Nguyen et al., 2016)), we use its training set to construct our demonstration pool $\mathcal{P}$. For each query in the training set, we construct positive and negative demonstrations by pairing the query with its relevant and irrelevant passages respectively. To maintain the output label balance in the demonstration pool $\mathcal{P}$, the number of negative demonstrations of each query is set equal to its positive demonstrations.
4.2. Demonstration Retriever DRetriever